A pendulum that has a period of 4.45 s and that is located where the acceleration due to gravity is 9.7 m/s2 is moved to a location where the acceleration due to gravity is 9.45 m/s2. What is its new period?

Round your answer to 5 decimal places


1
Expert's answer
2021-04-26T17:04:03-0400
"T=2\pi\sqrt{\frac{l}{g}}\\T'=\sqrt{\frac{g}{g'}}T\\T'=\sqrt{\frac{9.7}{9.45}}4.45=4.51\ s"